2/3=x/6 ?
x/2=16/6 ?
x/10=5/9 ?
9/4x=6/2 ?
4x/6=7/2 ?
a/5=3/2 ?
x-3/x=4/7 ?
y/y-16=5/3 ?

Answer by rfer(16322)

2/3=x/6
x=2/3*1/6
x=2/18
x=1/9
---------------
x/2=16/6
x=16/6*2/1
x=32/6
x=5 1/3